NEXT
Make New Entry,
Make Followup Entry
User name riordan
Log entry time 21:27:29 on March 03, 2009
Entry number 264397
keyword=Summary for March 2 Swing - March 3 Day
Summary for Mar 2 swing - Mar 3 Day
Awaiting recovery of machine until 12:00
Beam quality has improved significantly
Starting positron LHRS runs, doing Moller this evening
Running has been extended until 08:00, Monday March 16, 2009
Beam in Use
shift |
ABU |
BANU |
BNA |
2009/03/02 swing | 0.0 |
0.0 | 8.0 |
2009/03/03 owl | 0.0 |
0.0 | 8.0 |
2009/03/03 day | 2.4 |
1.0 | 4.6 |
total | 2.4 |
1.0 | 20.6 |
Main events/issues:
- Mar 2
- Shift on standby since no beam will be received
- Left Q3 current began fluctuating
- Mar 3
- Take controlled access to repair Q3. Found no apparent cause,
but after cleaning internal contacts the current seems stable.
- Beam is restored around 12:00. Positive polarity runs underway.
- BigBite deadtime seems to have gone up from 20% to 25% even
though rates are the same. Setting T2 prescale to 2 from 1.
- LHRS T3 trigger rate is 20x higher than negative polarity running.
Plans until day 3/5
- Moller, Evening March 3
- Continue positive polarity running until Day 3/4
- 06:30, go to controlled access for techs to swich LHRS polarity
- Finish as much 5 pass production as possible until beam studies
Tuesday in Kin I, G, ... with goals of 100% accumulation.
A copy of this log entry has been emailed to: brads, jpchen, jiang, meziani, choi, riordan